# X2I: Seamless Integration of Multimodal Understanding into Diffusion Transformer via Attention Distillation

### v2

@@ 4.3 Quantitative Results
IT2I Generation Tab. 2 shows the performance of different methods on the DreamBench. Overall, X2I has a slightly advantage in DINO scores, CLIP-I shows a two-point difference compared to KOSMOS-G and there is a considerable gap in CLIP-T compared to recent methods like OmniControl and IP-Adapter based on Flux.1, indicating that X2I still has room for improvement in its ability to faithfully follow the semantics of images.
@@ 4.4 Qualitative Results
X2I training achieves outstanding results through feeding pure text. Fig. 1 illustrates the effects of some of these functions. T2I possesses the ability to supports over 29 natural languages, covering the major language families worldwide; I2I includes image variations and the fusion of concepts from multiple images. IT2I encompasses single image instruction editing, style transfer, image stylization, instruction editing for multiple images, and the result of text token mixing with multiple images. It integrates multiple visual concepts to create unique images, extracts specific concepts from a set of input images, and synthesizes these concepts into new images, allowing precise control over the visual details of the images and avoiding common vagueness and uncertainty in text descriptions. Examples include changing clothes in the combination of clothes and people, performing group photos in person combinations, and creative conceptual combinations with three or more images.
@@
X2I enhanced Training. The purpose of constructing LightControl is to enhance the fidelity of image instruction editing. Essentially, it is not limited to a specific downstream task. In this paper, relevant experimental validation is only conducted in the stylization domain. As shown in Fig. 3, we simply chose the ink painting style for comparison. The first line represents the effect of DiffArtist, the middle highlighted in red is the reference image, and the second line is the effect of X2I enhanced Training. Compared to the DiffArtist method, X2I not only demonstrates a certain level of fidelity but also exhibits a strong ability to follow instructions. For more comparative results, please refer to the appendix Sec. 15.
Figure 3: The subjective results comparing the DiffArtist method
@@ 4.5 X2I Training Convergence
4.6 X2I’s reasoning ability and multi-turn dialogue generation ability
Figure 5: X2I’s reasoning ability.
Figure 6: X2I’s multi-turn dialogue generation ability.

### v3

@@ 4.3 Quantitative Results
IT2I Generation Tab. 2 shows the performance of different methods on the DreamBench. Overall, X2I has a slight advantage in DINO scores. CLIP-I shows a two-point difference compared to KOSMOS-G and there is a considerable gap in CLIP-T compared to recent methods like OmniControl and IP-Adapter based on Flux.1, indicating that X2I still has room for improvement in its ability to faithfully follow the semantics of images.
@@ 4.4 Qualitative Results
X2I training achieves outstanding results through feeding pure text. Fig. 1 illustrates the effects of some of these functions. T2I possesses the capability to support over 29 natural languages, covering the major language families worldwide; I2I includes image variations and the fusion of concepts from multiple images. IT2I encompasses single image instruction editing, style transfer, image stylization, instruction editing for multiple images, and the result of text token mixing with multiple images. It integrates multiple visual concepts to create unique images, extracts specific concepts from a set of input images, and synthesizes these concepts into new images, allowing precise control over the visual details of the images and avoiding common vagueness and uncertainty in text descriptions. Examples include changing clothes in the combination of clothes and people, performing group photos in person combinations, and creative conceptual combinations with three or more images.
@@
X2I enhanced Training. The purpose of constructing LightControl is to enhance the fidelity of image instruction editing. Essentially, it is not limited to a specific downstream task. In this paper, relevant experimental validation is only conducted in the stylization domain. For more comparative results, please refer to the appendix Sec. 16.
@@
Reasoning ability and multi-turn dialogue generation ability. MLLMs themselves possess certain reasoning abilities and multi-turn dialogue capabilities. X2I also inherits some of these related abilities. More results in the appendix Sec. 6.
@@ 5 Conclusion and Limitations
6 X2I’s reasoning ability and multi-turn dialogue generation ability
Figure 4: X2I’s reasoning ability.
Figure 5: X2I’s multi-turn dialogue generation ability.